Dear colleagues,
Here is a summary of today's activites at CSBF.
- Muon run in flight mode during overnight was running w/o stopping. Data copied to cdps3.
- Nick has looked at the muon data with the well tuned threshold values by Isaac and the TCD signals seem more reasonable even though MXR problem still prevent for the final conclusion.
- Chercam HSK display problem debugging continue. Brian and Jeremy confirmed that the connector was not modified during the test. Next step is to look at the HSK USB output directly.
- CD HV monitoring values calibrated again with the hsk display conversion as 1, which is raw voltage value, in the morning. And a set of formula obtained. The excel file was uploaded on the black board and the new conversion factors implemented into the hkconv4.dat. Now HV values seem consistent with the real values.
- A CD pulse shaper board delivered from PSU. Jason replaced the board and the CD signals from muon triggers look reasonable now.
- S3's bad smb connector was replaced with Jason's good one yesterday. The muon data taken with the new cable shows improvement even though still waiting for the MXR problem resolved for the final confirmation.
- Legs installed into the pallet and the CREAM-IV instrument moved next to the CDM for the integration.
- Power and Heater cables are routed and connected to the connectors.
- After the safe to mate test of power, sensor, and heater cables to the CDM, they are connected to the CDM.
- The temperature sensor cable for the CDM tested and found one connector for the CER -X-Y (J15) was connected to the wrong position. (J33) Also connectors for the J16 (SCD -Y PCB) and J18 (?), which are not necessary, were not available. Instead two sensor for SFC1 and SPAR. BOX STACK were made at UMD and connected to the two positions but no signals. Tried those sensors in different position and they were working. Problem with those positions? Need further investigation if we have time. Or just fly with existing sensors (16 out of 18 original) which are good enough.
- The SFC power on signal from the SFC interconnect module (Switch) did not work. It was fixed by removing pin #5 from the connector since both pin #4 and #5 had +4.7 V instead of just pin #4 thanks to Larry's suggestion. This bug fixed in the next version of the board but not with this board and Peter might did the same thing for the CREAM-III integration at WFF.
- Turned on the CREAM-IV instrument with the CDM power supply and muon run in flight mode with direct Ethernet connection between SFC and the CDM Client PC tested and OK.
- Ethernet cables between SFC1/2 and the CDM tried but not working. It turned out the connectors for the CAT5 cables were bad. The climped wires were loose. Will be fixed tomorrow with climping and soldering if necessary.
- Change of IP addresses of cdps1 and 2 found to be not possible in WFF side as we experienced before. Dr. Kim moved cdps1 and 2 to the lab for the remote commanding test on Monday and Tuesday. Still need to set the WFF IP addresses for unblocking.
- Second Ethernet card installation into the cdps5 tried but in vain by Prashant. We may have to buy a new card in the Radio shack and try it.
- Debugging of the reduced hsk data continued by Prashant.
- SangEun has been working on the SIP interface coding.
Here are things to do.
- Integration with the CDM continue with the Ethernet connection. The CDM module plan to be mounted under the instrument tomorrow.
- SFC Ethernet cable replacement with new ones. Existing ones are temporary with bad connectors.
- Reduced hsk debugging.
- Chercam HSK problem debugging.
- TCD 12 V HSK attenuation circuit installation.
- RPSC presentation preparation.
- Update of manuals including CPT.
- TCD S3 tuning.
- CD integration.
Please let us know if you have any questions/comments/suggestions.
Cheers, MooHyun.
MooHyun, July 18, 2008



